# The chromatic number of 4-dimensional lattices

### v2

@@ 1. Introduction
Equivalently, the chromatic number of a lattice $\Lambda$ is the chromatic number of the (infinite) Cayley graph on the additive group $\Lambda$ with strict Voronoi vectors as its set of generators. The strict Voronoi vectors of $\Lambda$, denoted by $\operatorname{Vor}(\Lambda)$, are the differences $v-w$ such that the intersection of the distinct translates $v+V(\Lambda)$ and $w+V(\Lambda)$ is a common facet. Thus,
@@
Many natural questions regarding $\chi(\Lambda)$ remain unanswered. For instance, it is not known whether $\chi(\Lambda)$ is computable. By the compactness theorem of de Bruijn and Erdős [3] we only know that the decision problem “Is $\chi(\Lambda)\leq k$?” is semidecidable [^1] There is an algorithm which terminates with a positive answer after finitely many steps on every input $\Lambda$ with $\chi(\Lambda)\leq k$. On inputs $\Lambda$ with $\chi(\Lambda)>k$, the algorithm either terminates with a negative answer or it runs in an infinite loop. because $\chi(\Lambda)$ is equal to the largest chromatic number of all finite, induced subgraphs of $\Lambda$. Directly related is the question of whether one can always find a periodic coloring with $\chi(\Lambda)$ colors. If so, and if we knew a priori a bound on the period, then this would immediately yield an algorithm for computing $\chi(\Lambda)$.
@@ 2. Voronoi graphs of lattices in four dimensions
Note that there is another possible starting point with the classification of iso-edge domains (also called $C$-type domains). For more information on the concept of iso-edge domains we refer to Ryshkov and Baranovskii [13] and to Dutour Sikirić and Kummer [9]. In general we now have $3$ equivalence relations regarding lattices, by identifying lattices according to either their Voronoi graph, their iso-edge domain, or their Delaunay subdivision. It is a priori clear that these relations become finer with Delaunay subdivisions being the finest one and Voroni graphs being the coarsest one. In dimension $4$ we have, up to equivalence, $51$ inequivalent iso-edge domains and $52$ inequivalent Delaunay subdivisions, so for our approach there is not much of a difference in using one of the two. We prefer to work with Delaunay subdivisions, as they are more common in the literature. Note that our classification of Voroni graphs in dimension $4$ yields that there are $16$ such, up to equivalence, so their number is strictly smaller then the number of iso-edge domains.
@@ 2.1. Classification of Delaunay triangulations
Here we recall some details of Voronoi’s classification of Delaunay triangulations of quaternary positive definite quadratic forms. We refer to the book by Schürmann [14] for a contemporary exhibition of Voronoi’s theory.
@@
At the end of his second memoir, Voronoi [17] computed the orbits of full-dimensional secondary cones for $n\leq 4$. For $n\leq 3$ there is only one orbit and for $n=4$ there are exactly three orbits.
@@ 16 $V_{8}^{z}$ $8$ $8$ $40320$ $1$ $1+1+1+1$
Here, we succeeded testing whether two Voronoi graphs are isomorphic in a rather ad hoc manner. This can also be done in a systematic way: In the next section we provide a finite algorithm that can be used in general, for this we will briefly discuss how to test for isomorphism of general Cayley graphs associated with lattices with finite sets of generators; see also the paper [1] by Baburin which summarizes computational approaches to finding isomorphisms and automorphisms of Cayley graphs of lattices and contains references to earlier works in this field.
@@ 3.2. The discrete torus bound
For any choice of a full-dimenssional $\Lambda^{\prime}$ we refer to the bound of the above lemma as the discrete torus bound induced by $\Lambda^{\prime}$.
@@ 3.3. Proof of the main theorem
The proof of our main theorem ultimately involves a finite list of computations. The general strategy is as follows. We separately consider each of the 16 isomorphism classes of Cayley graphs of four-dimensional lattices. For each such graph $C$, we compute the Delaunay polytope bound, that is the chromatic number of the subgraph $C_{1}$ (see Section 3.1). Next we select a lattice $\Lambda$ representing $C$ and choose a suitable sublattice $\Lambda^{\prime}$ such that the discrete torus bound induced by $\Lambda^{\prime}$ matches the Delaunay polytope bound. It turns out, simply by trial and error, that the sublattices $\Lambda^{\prime}$ can be chosen to be of the form $c\Lambda$ for some $c\in\mathbb{N}$, such that the bounds coincide. The information on which lattice and constant $c$ we used for our computation can be found in Table 3.1.
@@ 16 $V_{8}^{z}$ $1+1+1+1$ 2 2
We look for the smallest number $k$ such that there exists a coloring with $k$ color classes. This can be rephrased as a Boolean formula in conjunctive normal for (CNF) as follows. Write $V=\{v_{1},\ldots,v_{n}\}$; these are the variables. Assume that the colors are given by $1,\ldots,k$. We use a total of $n\cdot k$ variables
@@ Acknowledgments
The authors also express their gratitude to the suggestions of one of the anonymous referees. In comparison to an earlier draft of this manuscript we included a suggested strengthening of the statement of Lemmma 3.2, adapted question (2) from the referee report, and revised question (4) concerning periodic optimal colorings.

### v3

@@ 1. Introduction
Equivalently, the chromatic number of a lattice $\Lambda$ is the chromatic number of the (infinite) Cayley graph on the additive group $\Lambda$ with strict Voronoi vectors as its set of generators. The Voronoi vectors of $\Lambda$, are the differences $v-w$ such that the intersection of the distinct translates $v+V(\Lambda)$ and $w+V(\Lambda)$ is a common face; but not necessarily a common facet. A Voronoi vector as above is a strict Voronoi vector if this common intersection is a facet, we denote the set of such vectors by $\operatorname{Vor}(\Lambda)$. Thus,
@@
Many natural questions regarding $\chi(\Lambda)$ remain unanswered.
For instance, it is not known whether $\chi(\Lambda)$ is computable. By the compactness theorem of de Bruijn and Erdős [3] we only know that the decision problem “Is $\chi(\Lambda)\leq k$?” is semidecidable because $\chi(\Lambda)$ is equal to the largest chromatic number of all finite, induced subgraphs of $\Lambda$. This means that there is an algorithm which terminates with a positive answer after finitely many steps on every input $\Lambda$ with $\chi(\Lambda)\leq k$. On inputs $\Lambda$ with $\chi(\Lambda)>k$, the algorithm either terminates with a negative answer or it runs in an infinite loop.
Directly related is the question of whether one can always find a periodic coloring with $\chi(\Lambda)$ colors. If so, and if we knew a priori a bound on the period, then this would immediately yield an algorithm for computing $\chi(\Lambda)$.
@@ 2. Voronoi graphs of lattices in four dimensions
Note that there is another possible starting point with the classification of iso-edge domains (also called $C$-type domains). For more information on the concept of iso-edge domains we refer to Ryshkov and Baranovskii [14] and to Dutour Sikirić and Kummer [10]. In general we now have $3$ equivalence relations regarding lattices, by identifying lattices according to either their Voronoi graph, their iso-edge domain, or their Delaunay subdivision. It is a priori clear that these relations become finer with Delaunay subdivisions being the finest one and Voroni graphs being the coarsest one. In dimension $4$ we have, up to equivalence, $51$ inequivalent iso-edge domains and $52$ inequivalent Delaunay subdivisions, so for our approach there is not much of a difference in using one of the two. We prefer to work with Delaunay subdivisions, as they are more common in the literature. Note that our classification of Voroni graphs in dimension $4$ yields that there are $16$ such, up to equivalence, so their number is strictly smaller than the number of iso-edge domains.
@@ 2.1. Classification of Delaunay triangulations
Here we recall some details of Voronoi’s classification of Delaunay triangulations of quaternary positive definite quadratic forms. We refer to the book by Schürmann [15] for a contemporary exposition of Voronoi’s theory.
@@
At the end of his second memoir, Voronoi [18] computed the orbits of full-dimensional secondary cones for $n\leq 4$. For $n\leq 3$ there is only one orbit and for $n=4$ there are exactly three orbits. It turns out that in these cases, all full-dimensional secondary cones are simplicial. This is no longer the case for $n\geq 5$. For instance, in [7], for $n=6$, a full-dimensional secondary cone (of dimension $\binom{n(n+1)}{2}=21$) having $7145429$ many extreme rays was found.
@@ 16 $V_{8}^{z}$ $8$ $8$ $40320$ $1$ $1+1+1+1$
Here we were able to test whether two Voronoi graphs are isomorphic in a rather ad hoc way. This can also be done in a systematic way: In the next section we provide a finite algorithm that can be used in general, for this we will briefly discuss how to test for isomorphism of general Cayley graphs associated with lattices with finite sets of generators; see also the paper [1] by Baburin which summarizes computational approaches to finding isomorphisms and automorphisms of Cayley graphs of lattices and contains references to earlier works in this field.
@@ 3.2. The discrete torus bound
For any choice of a full-dimensional $\Lambda^{\prime}$ we refer to the bound of the above lemma as the discrete torus bound induced by $\Lambda^{\prime}$.
@@ 3.3. Proof of the main theorem
The proof of our main theorem ultimately involves a finite list of computations, although this was not guaranteed apriori. The general strategy is as follows. We separately consider each of the 16 isomorphism classes of Cayley graphs of four-dimensional lattices. For each such graph $C$, we compute the Delaunay polytope bound, that is the chromatic number of the subgraph $C_{1}$ (see Section 3.1). Next we select a lattice $\Lambda$ representing $C$ and choose a suitable sublattice $\Lambda^{\prime}$ such that the discrete torus bound induced by $\Lambda^{\prime}$ matches the Delaunay polytope bound. It turns out, simply by trial and error, that the sublattices $\Lambda^{\prime}$ can be chosen to be of the form $c\Lambda$ for some $c\in\mathbb{N}$, such that the bounds coincide. The information on which lattice and constant $c$ we used for our computation can be found in Table 3.1.
@@ 16 $V_{8}^{z}$ $1+1+1+1$ 2 2
We look for the smallest number $k$ such that there exists a coloring with $k$ color classes. This can be rephrased as a Boolean formula in conjunctive normal form (CNF) as follows. Write $V=\{v_{1},\ldots,v_{n}\}$; these are the variables. Assume that the colors are given by $1,\ldots,k$. We use a total of $n\cdot k$ variables
@@ Acknowledgments
The authors also express their gratitude to the suggestions of one of the anonymous referees. In comparison to an earlier draft of this manuscript we included a suggested strengthening of the statement of Lemma 3.2, adapted question (2) from the referee report, and revised question (4) concerning periodic optimal colorings.